OpenStudy (anonymous):

-5x-15=-32 solve step by step. why did I come up with an answer of -17/-5. and the correct answer is x=17/5

HanAkoSolo (jamierox4ev3r):

-5x-15=-32 +15 +15 ------------- -5x=-17 /-5 /-5 --------- x=\(\frac{17}{5}\)

HanAkoSolo (jamierox4ev3r):

hmmmm, idk if you can read that, but that is 17/5. You are correct, even tho it doesn't seem like it ^.^ good job :D

OpenStudy (anonymous):

I can read it. but why does it change negative to positive. 17/5

HanAkoSolo (jamierox4ev3r):

when both of the numbers are negative, they end up both becoming positive; that's just the way it is written :)

OpenStudy (uri):

Double negative =positive.

OpenStudy (anonymous):

ok thank you

HanAkoSolo (jamierox4ev3r):

right also 7-(-3)= 10. i think of it as the two negative signs can be turned into an addition sign, but -7-3=-10. Got it?

OpenStudy (anonymous):

check